OpenStudy (anonymous):

simplify the expression below and write it as a single logarithm: 3log(x+4)-2log(x-7)+5log(x-2)-log(x^2)

OpenStudy (luigi0210):

Do you know what to do with exponents?

OpenStudy (anonymous):

no could you please explain how to work it out please.

OpenStudy (luigi0210):

Well whenever there is a number infront of the log, you have to convert it to an exponent like such: \[2logx <=>logx^2\]

OpenStudy (anonymous):

oh okay

OpenStudy (luigi0210):

So can you convert the others?

OpenStudy (anonymous):

yea i think i can, thank you

OpenStudy (luigi0210):

Remember that: \[loga+logb=\log(a*b)\] \[loga-logb=\log \frac{a}{b}\]